OpenStudy (anonymous):

factor (p+3q)-r(3q+p)

terenzreignz (terenzreignz):

A little (just a little) tricky, but you only need to remember two things: FIRST a + b = b + a And SECOND ac + bc = c(a + b)

terenzreignz (terenzreignz):

So, the r(3q+p) may be written instead as r(p+3q) Turning this expression into (p+3q) - r(p+3q)

OpenStudy (anonymous):

so now i have (p+3q)-rp-3qr ?

terenzreignz (terenzreignz):

No. While that is technically correct, you're being asked to factor, not distribute. Distributing is in fact a step backwards when it comes to factoring... Just remember ac + bc ----> c(a + b)

OpenStudy (anonymous):

so what do i do ?

terenzreignz (terenzreignz):

Do you not see anything that looks like this? ac + bc ? Maybbe ac - bc ? It works on the same principle ac - bc = c(a - b)

terenzreignz (terenzreignz):

And you can be quite flexible when it comes to your choice of "c"

OpenStudy (anonymous):

can you tell me the setup ? then ill solve it

terenzreignz (terenzreignz):

Okay, why don't we let c = (p + 3q) So now, we have c - rc Does that look more agreeable to you? :)

OpenStudy (anonymous):

do you know what the answer is ?

terenzreignz (terenzreignz):

No ;)

OpenStudy (anonymous):

can you just set up the problem for me ?

terenzreignz (terenzreignz):

I am... We let c = (p + 3q) So that the expression becomes c - rc so use that second concept I told you, the one involving distribution...